Can someone please help me factorise the following
1) 18x^2-96x+128
2) 45y^4z-80y^2z^3
3) x^4+5x^3-7x^2-5x+6 given(x^2-1) is a factor.
Thank you very much

Found 2 solutions by stanbon, Vivek g:
Answer by stanbon(75887)

factorise the following
1) 18x^2-96x+128
2[9x^2-48x+64]
2(3x-8)(3x-8)
--------------------
2) 45y^4z-80y^2z^3
5y^2z[9y^2-16z^2]
5y^2z(3y-4z)(3y+4z)
-------------------------
3) x^4+5x^3-7x^2-5x+6 given(x^2-1) is a factor.
Using synthetic division:
---
1)....1....5....-7....-5....6
.......1....6.....-1....-6..|..0
-1)....1.....5.....-6....|..0
Quotient: x^2+5x-6
=======================
Factored Form: (x-1)^2(x+1)(x+6)
=====================================
Cheers,
Stan H.

Answer by Vivek g(3)

Age x yrs.
Age 30-x yrs.
X(30-x)=216
30x-x^2-216=0
x^2-30x+216=0. On sol. We get x=36. Ages are 6 and 36